Russell Preston delivers parts for several local auto parts stores. He charges clients $0.90 per mile driven. Russell has determined that if he drives 3,600 miles in a month, his average operating cost is $0.40 per mile. If he drives 5,600 miles in a month, his average operating cost is $0.35 per mile. Russell has used the high-low method to determine that his monthly cost equation is total cost = $640 + $0.26 per mile.

1. Break Even Miles

2. Assume Russell drove 1,500 miles last month. Without making any additional calculations, determine whether he earned a profit or a loss last month. Profit or Loss?

3. Determine how many miles Russell must drive to earn $800.00 in profit.

4A. Prepare a contribution margin income statement assuming Russell drove 1,500 miles last month.

4B. Use the information provided in Req 4a to calculate Russell’s degree of operating leverage. (Round your answer to the 2 decimal places.)
Answer #1

Contribution Margin per mile = Revenue per mile - variable cost per mile

= 0.90-0.26

= $0.64 per mile

1.Break even miles = Fixed costs/Contribution Margin per mile

= 640/0.64

= 1000 miles

2.Profit since miles driven are higher than break even miles

3.Number of miles = (Desired profit+Fixed costs)/Contribution Margin per mile

= (800+640)/0.64

= 2,250 miles

4.

Contribution Margin income statement
Sales revenue 1350
Less: Variable costs 390
Contribution Margin 960
Fixed costs 640
Operating Income 320
Degree of operating leverage = Contribution Margin/Net operating income 3 times